OpenStudy (anonymous):

solve by the elimination method 3x=42+2y 5x=4-4y

OpenStudy (anonymous):

all the examples my teacher gave did not involve a x on the other side

OpenStudy (haseeb96):

3x-2y=42......EQ 1 5x+4y=4.........EQ 2 mutiply eq 1 by 2 now it becomes 6x -4y =84 5x+4y=4 now eliminate 4y , it would be cancel 6x=84 5x=4 adding them 11x=88 x=88\11 x=8 now put the value of x in eq 2 5(2)+4y=4 4y=4-10 4y=-6 y=-6\4 y=-3\2 so the solution set is { ( 8,-3\2) } this is the correct answer

OpenStudy (anonymous):

I know but im also working on other problems because this work is timed i just wanted some input but thank you @Loser66